April 13

10:50 Opening the dome after waiting for a high wind to subside.

11:03 Starting with A851.

11:20 Having trouble identifying the field. Something's wrong with the acquisition script.

11:29 Switching to CLJ1226. Guide star guide#1 doesn't work.

11:37 Switching to guide star guide#2. Guider probe is failing to pick up the guide star (encoder problem).

12:10 Starting to acquire the source on the long-slit. Humidity's coming up (75%).

12:14 Some clouds are coming in. Lost the guide star.

12:22 Started the first science observation (long-slit spectroscopy of CLJ1226 in K), but clouds kept coming in. Guide star's lost... This exposure's no good. Script failed due to the loss of the guide star. Condensation appearing on the dewer window.

Stopped observing

12:53 Telescope put to the zenith. Trying to remove the condensation by increasing the air flow on the dewer window.

3:16 Clouds are gone but the seeing is blowing up to 3-4". Waiting for the condition to improve...

Resumed observing

5:10 Started K-band imaging observation of MACSJ1427 (backup imaging target for Greg). The seeing has improved to ~1.5-2".

5:23 Lost guiding because the sky got too bright. Stopped observing.

- End of night -

-- %USERSIG{Eiichi Egami - 2014-04-14}%

Comments

-- %BUBBLESIG{GregoryLeeWalth - 2014-04-14}%
 
Topic revision: r5 - 15 Apr 2014, GregoryLeeWalth
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback